Paul can do a job in two hrs..when Paul works alone ,he can do five jobs in 15 hrs.How many jobs can Mary do in 12 hrs, alone? Please show me step by step.
Answer by stanbon(75887)


Paul and Mary can do a job in two hrs..
when Paul works alone ,he can do five jobs in 15 hrs.
many jobs can Mary do in 12 hrs, alone? Please show me step by step.
----
Together rate: 1/2 job/hr
----
Paul rate: 1/3 job/hr (because he can do one of the jobs in 3 hrs)
----
Mary rate: 1/x job/hr (because she can do one of the jobs in x hrs)
======================
Equation:
rate + rate = together rate
1/3 + 1/x = 1/2
-----
2x + 6 = 3x
x = 6 hrs (time for Mary to do one of the jobs)
---
Ans: Mary can do 2 of the jobs in 12 hrs.
===============================================
cheers,
Stan H.